degree of operating leverage and financial leverage

reflects last year’s operations:

Sales $18,000,000
Variable costs 7,000,000
Revenue before fixed costs $11,000,000
Fixed costs 6,000,000
EBIT $5,000,000
Interest expense 1,750,000
Earnings before taxes (EBT) $3,250,000
Taxes 1,250,000
Net income $2,000,000

REQUIRED:
1. At this level of output, what is the degree of operating leverage?
2. What is the degree of financial leverage?
3. What is the degree of combined leverage?
4. If sales increase by 15%, by what percent would EBT (and net income) increase?
5. What is your firm’s break-even point in sales dollars?
Answer #1
1) Degree of operating leverage (DOL):
- It measures the EBIT's percentage change as a result of a change of
one percent in the level of output.
- It helps in measuring the business risk.
Degree of operating leverage = Sales revenue less total variable cost
divided by sales revenue less total cost:

DOL = (Sales-Variable Costs) / (Sales-Variable Costs-Fixed Costs)
ie DOL = $11,000,000/$5,000,000 = 2.2............Ans (1)

2) Degree of financial leverage.

The degree of financial leverage (DFL) is defined as the percentage
change in earnings per share [EPS] that results from a given
percentage change in earnings before interest and taxes (EBIT):
DFL = Percentage change in EPS divided by Percentage change in EBIT
ie DFL = EBIT / (EBIT-I), where I is the interest expense.
ie RFL = 5,000,000/$3,250,000 = 1.54..........................Ans (2)

3)Degree of Total leverage (DTL).
The degree of Total leverage is also known as degree of combined leverage
To compute it use the following formula:
DCL = DOL * DFL = 2.2*1.54 =3.39........................Ans (3)

4. If sales increase by 15%, by what percent would EBT (and net income) increase?
Sales 1.15*$18,000,000 = $20,700,000
Less Var costs 1.15* 7,000,000 = $8,050,000
---------------------------------------
Revenue before fixed costs $12,650,000
Less Fixed costs 6,000,000
---------------------------------------------
EBIT $6,650,000
Less Interest expense 1,750,000
----------------------------------------------------
Earnings before taxes (EBT) $4,900,000
Less Taxes 38.46% (1,884,540)
---------------------------------------
Net income $3,015,460

So EBT increase from $3,250,000 to $4,900,000
ie by ($4,900,000-3250,000)/3250,000 = 50.77% ..........Ans (4a)

Net Income increased from $2,000,000 to $3,015,460
ie by ($3,015,460-$2000,000)/$2000,000 = 50.77% ,.....Ans (4b)

5. What is your firm’s break-even point in sales dollars?
break-even point in sales dollars =company's total fixed expenses/contribution margin ratio.
CM Ratio = Contribution Margin/Sales = 11M/18M = 61.11%
So BEP $$$ = FC/CM Ratio = 6M/61.11% = $9.82M
BEP will not change with change in sales unles FC changes
answered by: nupur